OpenStudy (anonymous):

After travelling a certain distance, a train develops a snag and decreases its speed to half its original speed and reaches its destination 45 minutes late. Had the snag occurred 30 km further on, it would have reached its destination 15 minutes earlier. What is the speed of the train?

OpenStudy (anonymous):

put them in proportions!

OpenStudy (anonymous):

x-30/45+x/15=1

OpenStudy (anonymous):

x=75/2

OpenStudy (anonymous):

@vishweshshrimali5

OpenStudy (anonymous):

brb in 15 minutes.

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Let 's' be the speed of the train and 'x' be the distance where the snag occurred. 

OpenStudy (anonymous):

yes how to frame the equation.

OpenStudy (aum):



OpenStudy (anonymous):

s/2*45=x,s/2*(x+30)(45-30)=x

OpenStudy (aum):

Assume it takes 'T' minutes to reach the destination if there are no snag. speed = distance / time. time = distance / speed From the top part of the diagram we have: x / s + y / (s/2) = T + 45 x/s + 2y/s = T + 45 ------ (1) From the bottom part of the diagram we have: (x+30) / s + (y-30)/ (s/2) = T - 15 (x+30) / s + 2(y-30)/ s = T - 15 ------ (2) Subtract (2) from (1): x/s + 2y/s - x/s - 30/s - 2y/s + 60/s = 60 30/s = 60 s = 30/60 = 1/2 km / minute s = 1/2 * 60 = 30 km / hr.

OpenStudy (anonymous):

how? time /speed=y/s/2

OpenStudy (anonymous):

something went wrong .i think

OpenStudy (aum):

time = distance / speed It travels distance 'y' at speed 's/2' time = y divided by s/2

OpenStudy (anonymous):

distance/speed=time.right? u have taken y time as distance,

OpenStudy (aum):

y is not time. 'y' is the distance from where the snag happens to the destination.

OpenStudy (aum):

'x' is the distance from the starting point to where the snag happens. 'y' is the distance from where the snag happens to the destination.

OpenStudy (anonymous):

but the answer option,90,100,105,120km/hr

OpenStudy (anonymous):

brb

OpenStudy (aum):

I think it is because of the poor wording of the problem: "Had the snag occurred 30 km further on, it would have reached its destination 15 minutes EARLIER." By 15 minutes earlier they don't mean 15 means earlier than normal time but 15 minutes earlier compared to the previous late time! In the first scenario the train arrived 45 minutes late. In the second scenario they are saying the train arrived 15 minutes earlier compared to the 45 minutes late which means the train arrived 30 minutes late instead of 45 minutes late. So in equation (2), on the right hand side, we need to change it to T + 30. x/s + 2y/s = T + 45 ------ (1) (x+30) / s + 2(y-30)/ s = T + 30 ------ (2) subtract x/s + 2y/s - x/s - 30/s - 2y/s + 60/s = 15 30/s = 15 s = 30/15 = 2 Km / min s = 60 * 2 = 120 Km / hr.

OpenStudy (aum):

In the first scenario the train arrived 45 minutes late. In the second scenario the train arrived 30 minutes late. And this what they mean by the train arriving 15 minutes earlier.